# Boosting Large-scale Parallel Training Efficiency with C4 : A Communication-Driven Approach

### v1

@@
Abstract.
The emergence of Large Language Models (LLMs) has necessitated the adoption of parallel training techniques, involving the deployment of thousands of GPUs to train a single model. Unfortunately, we have found that the efficiency of current parallel training is often suboptimal, largely due to the following two main issues. Firstly, hardware failures are inevitable, leading to interruptions in the training tasks. The inability to quickly identify the faulty components results in a substantial waste of GPU resources. Secondly, since GPUs must wait for parameter synchronization to complete before proceeding to the next round of computation, network congestions can greatly increase the waiting time for GPUs. To address these challenges, this paper introduces a communication-driven solution, namely the C4 . The key insights of C4 are two folds. First, in parallel training, collective communication exhibits periodic and homogeneous characteristics, so any anomalies are certainly due to some form of hardware malfunction. By leveraging this feature, C4 can rapidly identify the faulty components, swiftly isolate the anomaly, and restart the task, thereby avoiding resource wastage caused by delays in anomaly detection. Second, the predictable communication model of collective communication, involving few large flows, allows C4 to efficiently execute traffic planning, substantially reducing network congestion. C4 has been extensively implemented across our production systems, cutting error-induced overhead by roughly 30% and enhancing runtime performance by about 15% for certain applications with moderate communication costs.
1. Introduction
@@
The rapid advancement of Large Language Models (LLMs) (Radford et al., 2018, 2019; Brown et al., 2020; Touvron et al., 2023; Du et al., 2022; Bai et al., 2023) has significantly impacted the fields of machine learning and artificial intelligence. LLMs such as GPT-3 (Radford et al., 2018) and Llama (Touvron et al., 2023) are exhibiting human-like proficiency in text generation, question-answering, and even code creation. Despite the promising capabilities of these models, training them to reach their full potential is a complicated and resource-intensive process. Training a GPT-like model that has 175 billion parameters, for example, may require up to two months in a cluster of 1000 GPUs.
@@
A key challenge in training LLMs in production AI clusters is that the effective utilization of GPU is relatively low, due to the following two reasons. First, a substantial amount of time is consumed in diagnosing system errors throughout the training lifespan due to the high failure rate of cutting-edge GPU products (NVIDIA, [n. d.]a, [n. d.]c). Prolonged training sessions necessitate a stable process; however, the latest generation GPUs tend to exhibit high error rates, likely a consequence of their rapid development, rushed delivery, and increased power consumption. When running in Bulk Synchronous Parallel (BSP) (Valiant, 1990) mode, errors in any node can cause the entire job to fail. This necessitates extensive time for system diagnosis and node isolation before we restart the job.
@@
Furthermore, tuning the training process of these models to peak efficiency is a complex task, with communication costs being a significant impediment to scalability. That is, GPUs may experience delays at synchronization points while awaiting the outcomes of collective operations. The state of the arts suggest managing network traffic to improve communication performance. However, a shared physical cluster with multiple tenants running concurrent training jobs adds to the complexity of these patterns, exacerbating the difficulty in effective management. Additionally, the appearance of downlinks, which is increasingly common in large-scales systems, also diminishes the efficacy of previous traffic engineering strategies.
Figure 1. Problems that may occur in the lifetime of AI training applications.
This paper proposes C4 (Calibrating Collective Communication over Converged Ethernet), an innovative system designed to unleash the computational capabilities of massive collaborating GPUs in large-scale AI clusters. C4 is composed of two subsystems, including C4D (C4 Diagnosis) and C4P (C4 Performance). C4D aims to enhance training stability by reducing stalls caused by uncorrectable errors. C4D achieves this by automatically detecting system errors in real-time, isolating faulty nodes, and facilitating application restarts from the last checkpoint. C4D’s innovation includes the following designs. (1) We enhance the collective communication library to offer the capability to monitor its status when running collective operations. (2) By collecting and analyzing the monitored status from different workers, we could detect the potential errors when one or some of ranks present suspected syndromes. (3) Once a suspect rank/node is reported, we could automatically isolate it from the scheduling system, and restart the job from last checkpoint. (4) C4D also collects the data from other system monitors, including the job monitor, server monitor, and network monitor, and conduct offline analysis accordingly.
C4P is designed to reduce the collective communication cost within large-scale training clusters. It accomplishes this by uniformly distributing network connections across all available paths and dynamically adjusting the load on each path based on the real-time network conditions. The key challenges it addresses include managing concurrent jobs and handling link errors. The novelty of C4P lies in the following aspects: (1) The communication library is enhanced with the ability to request path allocation prior to establishing connections. This feature enables it to control the designated communication paths for each link and to dynamically redistribute the network load across these paths. (2) C4P incorporates a mechanism for monitoring the integrity of network links, allowing it to identify and exclude faulty links from being considered in future path allocations, thereby maintaining network health and reliability. (3)For every communication request, C4P selects an optimal path, taking into account the current load distribution across all available network paths, ensuring efficient and balanced utilization of network resources.
@@
C4 has been deployed in our AI training clusters for six months, and is now serving over 20 customers for LLM training. Leveraging its automatic error detection and recovery system, C4 effectively removes the majority of error-induced overhead, typically accounting for about 30% of the total time. Additionally, C4 mitigates the communication costs associated with imbalanced traffic. For the tested applications with modest communication cost, the throughput of jobs can be improved by approximately 15%. Collectively, these improvements result in an overall boost in cluster compute efficiency from 30% to 45%.
2. Challenges and Opportunities
@@
Our team has successfully deployed numerous AI clusters, equipped with hundreds of thousands of GPUs, to enable a wide range of concurrent AI training jobs. In contrast to conventional cloud services, we have observed a pronounced difference in the expected delivery standards. That is, the users of AI clusters demand that their tasks perform at peak efficiency due to the following two reasons. First, the value of the developed models is intrinsically tied to how swiftly they can be brought to market, making expedited training a priority. Second, the significant capital invested in the training infrastructure means that any performance degradation translates directly into substantial financial repercussions. However, maintaining a consistently perfect operating state for these large-scale systems proves challenging over extended periods, especially when the number of GPUs in a single job grow to thousands, or even tens of thousands.
The key challenges that forbid distributed training system operate in their peak performance include two aspects, as shown in Fig. 1. Firstly, the GPU worker may not be in the running state because of the occurrence of errors. Moreover, the GPU workers could wait for the completion of gradients synchronization and waste their computing power, even in running state.
Challenges and Opportunities in Error Handling. The distributed training jobs are sensitive to errors, because of the synchrony of collective communication. That is, any corrupted worker could cause the jobs to fail. To keep the training job moving forward, the users save checkpoints in run-time. And it will fallback to last valid checkpoint, if the job fails by some reasons. As shown in Fig. 1, critical errors can arise before or during the execution of DL training jobs. And they will affect the system utilization in different ways. If the defects exists before the start of training jobs, they will incur extra time spending on system diagnosing, error component isolation, and job restart. In this case, a lot of time is spent on system diagnosing and pinpointing the defect components, which may take hours or even days. While if the uncorrectable errors are encountered during the execution of training jobs, some more extra time is wasted, including the post-checkpoint cost and fault detection cost. Since the job must revert to the last valid checkpoint following an uncorrectable fault, the execution time after the last checkpoint is rendered futile. We refer to this as the post-checkpoint cost. Additionally, there is a delay between the occurrence of an error and its detection by users, which we term the fault detection cost.
Tab. 1 shows the statistics summary of errors encountered over a month by a representative job six months ago. The data indicates that the job experienced 40 crashes within that month due to these errors. In lack of efficient fault detection and diagnosing tools at that time, it can take hours to days to identify the root causes and pinpoint the defect nodes. Based on our experience, approximately 30% of time was expended on error detection, system diagnosis, isolating defective nodes, and restarting processes, which left only 70% of the time available for actual computing tasks.
Users typically perceive most error types as a generic “NCCL Error”, making it challenging to pinpoint the exact root causes and respond to the reported errors effectively. For example, hardware issues in a GPU might cause one of the worker processes to crash. As a result, its peer workers fail to receive an acknowledgment after sending a message to the crashed worker, leading to an error code of 12 being thrown by the communication library. This same error code could also arise from other issues such as network disconnections, memory overflows causing container eviction, or even user code errors, such as tensor size mismatches. A key insight from the table is that the majority of errors (about 82.5%), are confined to specific nodes or even individual devices, suggesting the possibility of isolating the localized faulty components. This isolation strategy could prevent subsequent jobs from being impacted by the same faults.
Table 1. Errors classification in AI training clusters.
Users’ View Root Causes Error Rate Localized
@@
To minimize recovery costs, it is essential to address each contributing factor appropriately. For swift diagnosis, our system automatically identifies defective nodes by leveraging the inherent characteristics of deep learning training jobs, such as balanced computing costs within a data parallel group and synchronization between iterations. Upon locating a defect, the system promptly initiates isolation and restart procedures, deferring in-depth root cause analysis to offline processing. To decrease the time spent on fault detection, human intervention should be eliminated, expediting the process from tens of minutes to mere seconds. In addition, to curtail the post-checkpoint cost, we have implemented a rapid checkpointing mechanism, similar to the prior work (Wang et al., 2023a), capable of saving checkpoints approximately every 10 iterations.
Figure 2. Scalability loss grows with the scale of systems.
@@
Challenges and Opportunities in Optimizing Communication Performance. In addition to uncorrectable faults leading to job crashes, GPUs can experience stalls due to delays in completing collective operations or data loading tasks. This paper concentrates on addressing unnecessary overhead in collective operations. To begin with, elevated communication costs may arise from traffic collision among various flows or hardware issues, such as degraded PCIe lanes or Ethernet link failures. To simultaneously expand the maximum system scale and improve system availability and maintainability, we have adopted dual-port Network Interface Cards (NICs) for system interconnects, with two separate ports bonded together. Having each port connecting to a distinct leaf switch, there would be two available links between the hosts and leaf switches. Should one link become inoperative, the remaining link would take over all the downstream traffic, making it a bottleneck in the system. Furthermore, beyond the communication overhead caused by hardware defects, there is also the potential for multiple data flows to compete for the available bandwidth of a single link. For today’s distributed training jobs, bandwidth constraints can significantly prolong the communication latency and lead to degradation of communication performance. As modern models and the training clusters grow in scale, this problem would be further exacerbated. Fig. 2 presents a comparison of actual versus ideal performance, training a representative GPT-like model composed of 22 billion parameters. It is evident that there exists a discrepancy between the effective performance achieved and the theoretical ideal performance, and this disparity amplifies as the scale of the system increases. Notably, when the system scales to 512 GPUs, the effective performance drops to 30% below the ideal performance.
@@
Our analysis reveals that communication patterns in DL training clusters markedly differ from those in traditional cloud environments. In conventional cloud settings, we commonly observe a large number of concurrent connections, often several tens of thousands per instance. Such a volume of connections can be distributed across the network with relatively low variation using the Equal-Cost Multi-Path (ECMP) (Hopps, 2000) routing strategy. Conversely, in distributed training systems, there are only a few longer-lasting connections, with each node typically managing around a few hundred connections. Traffic collisions in this environment can lead to substantial variations in the effective bandwidth of these connections, resulting in multi-fold increases in communication delays. However, this challenge is accompanied by an opportunity: the limited number of connections and their predictable patterns in DL training clusters offer a unique advantage for enhancing communication efficiency through comprehensive global traffic engineering.
@@
In essence, the effective performance of large-scale training clusters is significantly influenced by hardware defects and traffic collisions. To ensure that GPUs operate efficiently during model computation, it is critical to minimize the time lost in fault detection and system diagnosis. Furthermore, to prevent GPU stalls during periodic synchronization, it is essential to eliminate any superfluous communication overhead.
3. Design

### v2

# Enhancing Large-Scale AI Training Efficiency: The C4 Solution for Real-Time Anomaly Detection and Communication Optimization

@@
Abstract
The emergence of Large Language Models (LLMs) has necessitated the adoption of distributed training techniques, involving the deployment of thousands of GPUs to train a single model. Unfortunately, the efficiency of large-scale distributed training systems is often suboptimal due to the increased likelihood of hardware errors in high-end GPU products and the heightened risk of network traffic collisions. Specifically, GPUs involved in the same job require periodic synchronization to exchange necessary data, such as gradients, parameters, or activations. As a result, any local hardware failure can disrupt training tasks, and the inability to swiftly identify faulty components leads to a significant waste of GPU resources. Moreover, prolonged communication due to traffic collisions can substantially increase GPU waiting times.
To address these challenges, we propose a communication-driven solution, namely the C4. The key insights of C4 are twofold. First, the load in distributed training exhibits homogeneous characteristics and is divided into iterations through periodic synchronization, therefore hardware anomalies would incur certain syndrome in collective communication. By leveraging this feature, C4 can rapidly identify the faulty components, swiftly isolate the anomaly, and restart the task, thereby avoiding resource wastage caused by delays in anomaly detection. Second, the predictable communication model of collective communication, involving a limited number of long-lived flows, allows C4 to efficiently execute traffic planning, substantially reducing bandwidth competition among these flows. The C4 has been extensively deployed across real-world production systems in a hyperscale cloud provider, yielding a significant improvement in system efficiency, from 30% to 45%. This enhancement is attributed to a 30% reduction in error-induced overhead and a 15% reduction in communication costs.
I Introduction
@@
The rapid advancement of Large Language Models (LLMs) [43, 44, 5, 51, 15, 2] has significantly impacted the fields of machine learning (ML) and artificial intelligence (AI). LLMs such as GPT [43] and Llama [51] families are exhibiting human-like proficiency in text generation, question-answering, and even code creation. Despite the promising capabilities of these models, training them to reach their full potential is a complicated and resource-intensive process. Training a GPT-like model that has 175 billion parameters, for example, may require up to two months in a cluster of 1000 GPUs [33].
@@
One of the major hurdles we faced when training LLMs is the suboptimal utilization of GPU resources, which can be attributed to two primary reasons. First, a significant amount of time (over 30%) is spent on recovering from system errors throughout the training lifespan. Extended training sessions necessitate a stable process; however, the latest generation GPUs [35, 37] tend to exhibit high error rates, likely a consequence of their rapid development, rushed delivery, and increased power consumption. When running in Bulk Synchronous Parallel (BSP) [52] mode, errors in any node can cause the entire job to fail. This necessitates extensive time for system diagnosis and node isolation before we restart the job.
@@
Furthermore, tuning the training process of these models to peak efficiency is a complex task, with communication costs being a significant impediment to scalability. That is, GPUs may experience delays at synchronization points while awaiting the outcomes of collective operations. The state of the arts [13, 14] suggest managing network traffic to improve communication performance. However, a shared physical cluster with multiple tenants running concurrent training jobs adds to the complexity of these patterns, exacerbating the difficulty in effective management. Additionally, the appearance of downlinks, which is increasingly common in large-scales systems, also diminishes the efficacy of previous traffic engineering strategies.
In order to tackle these challenges, this paper introduces Calibrating Collective Communication over Converged Ethernet (C4), an innovative system designed to unleash the computational capabilities of massive collaborating GPUs in large-scale AI clusters, by harnessing the inherent characteristics of distributed training. Specifically, it minimizes downtime caused by hardware errors by enabling the design of a highly efficient error recovery mechanism. Leveraging the homogeneous running rhythm of the distributed workers, it enables real-time fault detection by analyzing the timing discrepancies among them at the synchronization points, namely, the collective communication operations in each step. Moreover, C4 improves runtime communication efficiency through global traffic engineering. By exploiting the features of network traffic in AI clusters, C4 implements a cluster-level path management mechanism, mitigating the impact of traffic collisions and ensuring seamless collaboration among GPUs.
The contributions of this paper include:
  • We extend the capabilities of Alibaba Collective Communication Library (ACCL) [14] to support real-time status monitoring and communication path control, enabling the development of real-time error detection and global traffic engineering.
  • We design and implement a real-time hardware error detection subsystem that collects and analyzes the runtime status of ACCL across different workers, and builds upon this a fast recovery system that automatically isolates faulty nodes and restarts the affected jobs.
  • We develop a global traffic engineering subsystem that manages all the networking path resources in a cluster, and intelligently controls path allocation for establishing connections in ACCL, ensuring balanced load distribution across available paths.
@@
  • We successfully deploy C4 in our AI clusters, serving key customers for LLM training. It effectively removes the majority of error-induced overhead, typically accounting for about 30% of the total time. Additionally, C4 mitigates the communication costs, and improves the system throughput by approximately 15%.
II Understanding the Challenges in Operational AI Clusters
II-A New Challenges
In response to the immense success of LLMs, cloud providers have deployed dedicated AI clusters to effectively manage the growing demand for large-scale training tasks. Unlike traditional data centers for traditional cloud applications, large-scale AI clusters face distinct challenges arising from emerging workloads and rapidly evolving hardware systems.
One of the key challenges is that distributed training jobs are highly sensitive to local failures, which occur with notable frequency in modern large-scale training systems. This sensitivity arises from the periodical synchronization required among workers in each training step. It is well-known that contemporary LLMs are capable of harnessing tens of thousands of GPUs for their distributed training processes. The mainstream distributed training frameworks [49, 45] support a variety of parallel training strategies across multiple dimensions. With these parallel training strategies, each GPU manages a fraction of the model or input for localized computations. The workers within the same parallelism group must exchange gradients, parameters, or activations for further processing. If an exception occurs in any worker, it will block subsequent workers and cause the entire job to crash. Consequently, distributed training jobs, similar to HPC applications, must frequently save checkpoints and restart the entire job from the last valid checkpoint in the event of errors. What’s worse, our experience has shown that high-end GPU products tend to have higher error rates, increasing the probability a distributed training job may crash, especially as the system scales out. The recovery from job crashes can be costly and time-consuming, leading to a significant waste of computational capacity.
@@ II-A New Challenges
Another key challenge comes from the pronounced disparity in expected delivery standards of AI clusters. Specifically, users demand peak task performance, driven by two primary motivations. First, the value of the developed models is intrinsically tied to how swiftly they can be brought to market, making expedited training a priority. Second, the significant capital invested in the training infrastructure means that any performance degradation translates directly into substantial financial repercussions. However, maintaining a consistently optimal operational state for these large-scale systems is particularly challenging over extended periods, especially as the number of GPUs in a single job reaches thousands or even tens of thousands.
II-B Pain Points and Their Influences on Training Jobs
Fig. 1 illustrates the various issues that can negatively impact system utilization, including 1 anomalies in computing nodes, 2 failures and 3 traffic collision in networking subsystem, 4 I/O hang or slow in storage subsystem, and 5 problems in users’ code or supporting software. First of all, nodes with critical failures may cause job crashes. And some of the computing nodes could exhibit low performance, commonly referred to as slow nodes, often due to defects in GPUs or NICs. The likelihood of encountering defect nodes increases with system scales, ultimately undermining the efficiency of the entire system. Furthermore, traffic collisions among the workers, whether from the same or different jobs, can prolong the communication process and subsequently reduce system throughput. Link and switch failures in network subsystem could exacerbate the extent of traffic collision. In addition, issues in storage subsystem and software layer would systematically affect all the workers.
Figure 1: Issues that can adversely affect system utilization.
Figure 2: The influences of system issues on training jobs: 1) job failures (at startup and runtime) incurred overhead, and 2) slowdowns in communication or non-communication processes (computing or data loading).
From users’ perspective, system utilization is impaired in two distinct ways, as illustrated in Fig. 2. First, critical failures cause frequent job crashes, leading to a waste of computing power. Failures encountered during system startup necessitate additional time for system diagnosis, error component isolation, and job restart. While, if failures occur at runtime, there are additional overheads including post-checkpoint and fault detection costs. Upon crashing, the job must revert to the last valid checkpoint, rendering the execution time after the last checkpoint useless, which we refer to as post-checkpoint cost. There is also a delay between the occurrence of an error and the users’ awareness, known as the fault detection cost. Besides job crashes, non-critical issues at runtime, such as performance degradation in GPUs or PCIe fabric, and network traffic collisions, can also adversely affect system throughput. The extended communication time is referred to as communication slowdown, while the prolonged computing or data loading time is referred to as non-communication slowdown. Any overhead in communication or non-communication indicates degradation in system throughput. We would discuss the causes and effects of job crashes and runtime slowdowns respectively.
II-C The Quantitative Analyses of Job Crashes.
To investigate the underlying causes of job crashes and how the system utilization is impacted, we chose a representative job (one of jobs with largest scale) for analysis. This particular job, running in a newly deployed cluster, witnessed a total of 40 crashes within one month. We categories these crash causes and show their distribution in Tab. I.
The table reveals that it’s challenging to discern the underlying causes from the user’s perspective, as the majority of them simply show themselves as “NCCL Errors”. This is largely due to the synchronous nature of distributed training tasks. For instance, a hardware issue in a GPU could lead to the crash of one of the worker processes, causing its peer workers to fail to receive an acknowledgment after sending a message to the crashed worker. Consequently, this triggers an error code of 12 in the communication library. This same error code could also be generated by other issues such as network disconnections, memory overflows resulting in container eviction, or users’ code errors like tensor size mismatches. Regrettably, the industry lacks efficient tools to identify the root causes of these errors in a timely manner.
TABLE I: The distribution of crash causes recorded over a one-month from a representative job employing 4,096 GPUs.
Users’ View Root Causes Proportion Local
@@ II-B Pain Points and Their Influences on Training Jobs
This table also reveals an important fact: the majority of errors (approximately 82.5%) are confined to specific nodes or individual devices, largely attributed to the elevated error rates in high-end GPU productions. This indicates the potential for isolating the localized faulty components, presenting an opportunity to tackle the majority of errors [^1] Other systematic faults, such as faults in distributed storage subsystem, would be handled by dedicated monitoring tools. Nevertheless, due to the similarity in symptoms for most failures, manually identifying the root causes and locating the defective nodes could consume hours to days, particularly in a cluster with thousands of GPUs.
Figure 3: Performance loss grows with system scale, due to the increased extent of traffic collision.
@@
In our experience, around 30% of our time was dedicated to error detection, system diagnosis, isolating defective nodes, and restarting processes, leaving only 70% for actual computing tasks, in the absence of efficient fault detection and diagnostic tools. To minimize the failure-incurred costs, it is essential to address each contributing factor appropriately. For swift diagnosis, our system automatically identifies defective nodes by leveraging the inherent characteristics of training jobs, such as balanced computing costs within a data parallel group and synchronization between iterations. Upon locating a defect, the system promptly initiates isolation and restart procedures, deferring in-depth root cause analysis to offline processing. By developing automatic fault detection tool, we can eliminated human intervention, expediting the process from tens of minutes to mere seconds. In addition, to curtail the post-checkpoint cost, we have implemented a rapid checkpoint mechanism, similar to the prior work [53], capable of saving checkpoints approximately every 10 iterations.
II-D The Quantitative Analyses of Runtime Slowdowns.
Since non-critical failures in individual nodes can be addressed using the same strategy as critical errors, we will put emphasis on communication slowdowns in network. To begin with, we designed a specialized high-performance network architecture for AI clusters, with each computing node equipped with multiple high-performance Network Interface Cards (NICs) supporting Remote Direct Memory Access (RDMA) to facilitate fast inter-node communication among the heavily packed GPUs in each node. Each NIC provides two physical ports, connected to a pair of leaf switches, which are interconnected to spine switches through the Fat-Tree topology, with a 1:1 oversubscription ratio. This architecture offers distinct advantages: It provides high system availability and maintainability, allowing users to react more flexibly to non-critical link failures, such as scheduling timely restarts or initiating on-demand checkpoints that are strategically based on recent checkpointing activities. Additionally, it expands the maximum system scale by doubling the number of spine switches in the system, consequently doubling the total computational capacity of GPUs.
@@
Just as a coin has two sides, this architecture also presents challenges in communication performance. It’s known that elevated communication costs may arise from traffic collision among various flows. When numerous flows contend for the limited bandwidth of a single link, the flow completion time can increase by several times. In collective communication, any flow that is throttled can have a ripple effect, hindering the entire communication group from proceeding with subsequent operations. Moreover, a new problem arises with dual-port NICs. If there is an imbalance in the loads on the two physical ports, the port with the higher load can become a bottleneck. For today’s distributed training jobs, these issues can significantly prolong the communication latency and lead to degradation of communication performance. As modern models and the training clusters grow in scale, this problem would be further exacerbated. Fig. 3 presents a comparison of actual versus ideal performance, training a representative GPT-like model composed of 22 billion parameters. It is evident that there exists a discrepancy between the effective performance achieved and the theoretical ideal performance, and this disparity amplifies as the scale of the system increases. Notably, when the system scales to 512 GPUs, the effective performance drops to 30% below the ideal performance.
@@
Our analysis reveals that communication patterns in AI clusters markedly differ from those in traditional cloud environments. In conventional cloud settings, we commonly observe a large number of concurrent connections, often several tens of thousands per instance. Such a volume of connections can be distributed across the network with relatively low variation using the Equal-Cost Multi-Path (ECMP) [19] routing strategy. Conversely, in distributed training systems, there are only a few longer-lasting connections, with each node typically managing around a few hundred connections. Traffic collisions in this environment can lead to substantial variations in the effective bandwidth of these connections, resulting in multi-fold increases in communication delays. However, this challenge is accompanied by an opportunity: the limited number of connections and their predictable patterns in AI clusters offer a unique advantage for enhancing communication efficiency through comprehensive global traffic engineering.
@@
In a nutshell, the effective performance of large-scale training clusters is significantly influenced by hardware defects and traffic collisions. To maximize GPU utilization, it is crucial to reduce the time spent on fault detection and system diagnosis, thereby minimizing downtime. Additionally, eliminating unnecessary communication overhead is vital to prevent GPU stalls during periodic synchronization, ensuring optimal runtime performance.
III Design